OpenStudy (anonymous):

Suppose f (x) = x3 + x and g(x) = f -1(x). Find g'(2).

OpenStudy (anonymous):

what is the meaning of g(x)=f-1(x) \[g(x)=\frac{ 1 }{ f(x) }\]?

OpenStudy (anonymous):

inverse of f(x)

OpenStudy (anonymous):

\[g(f(x)=x\\ g'(f(x))f'(x)=1\\ g'(f(1) f'(1) =1\\ g'(2)=\frac 1 {f'(1)}\] Can you finish it?

OpenStudy (anonymous):

\[ f(x)= x^3 + x\\ f'(x) = 3 x^2 +1\\ f'(1) =4\\ g'(2)=\frac 1 4 \]